OpenStudy (anonymous):

Linear Algebra: Find conditions on a, b, and c such that the system has (i) no solution, (ii) infinitely many solutions, (iii) a unique solution.

OpenStudy (anonymous):

\[3x + y - z = a\] \[x - y + 2z = b\] \[5x + 3y - z = c\]

OpenStudy (anonymous):